# Odd-One-Out: Anomaly Detection by Comparing with Neighbors

### v2

@@ Abstract
This paper introduces a novel anomaly detection (AD) problem that focuses on identifying ‘odd-looking’ objects relative to the other instances in a given scene. In contrast to the traditional AD benchmarks, anomalies in our task are scene-specific, defined by the regular instances that make up the majority. Since object instances may be only partly visible from a single viewpoint, our setting employs multiple views of each scene as input. To provide a testbed for future research in this task, we introduce two benchmarks, ToysAD-8K and PartsAD-15K. We propose a novel method that constructs 3D object-centric representations from multiple 2D views for each instance and detects the anomalous ones through a cross-instance comparison. We rigorously analyze our method quantitatively and qualitatively on the presented benchmarks. The datasets, source code, and models will be made available upon publication.
@@ 1 Introduction
Anomaly detection (AD) [10, 29] aims to detect patterns that deviate from expected behavior. In standard AD benchmark for computer vision, anomalies may arise from high-level factors such as introduction of an object instance from an unseen category [9, 31, 2] or low-level ones such as defects and variations in the normal object shape and texture [8, 6, 13]. In these benchmarks, AD methods analyze each object individually, ignoring relationships between objects in the same scene or contextual factors. However, in many real-world scenarios, ‘normality’ can be conditional. For instance, in a production line for blue-handled coffee cups, a yellow handle is anomalous, and in a production line for yellow-handled coffee cups, a blue handle is anomalous (see Fig. 1(b)). This context dependent normality cannot be addressed by methods assuming a fixed standard of normality.
Inspired by real-world visual inspection, we introduce a new AD problem with two new benchmarks and a novel solution. In our setup (see Fig. 1(a)), each image contains a scene with multiple instances of the same object (e.g., coffee cup) including either all normal, or a mix of normal and anomalous instances. Since defining exact shape and appearance specifications on paper is often impractical, we assume each scene has multiple normal instances together providing a scene-specific reference for ‘normality’. Our objective is to detect anomalous instances in a scene and to generalize to previously unseen scenes including unseen objects and layouts. Unlike existing benchmarks, this task often requires a holistic understanding of the scene by comparing instances with each other with the exception of few anomalies such as cracks and misaligned parts that can be identified by inspecting individual instances. Additionally, since performing AD from a single view can be ambiguous due to self-occlusion and occlusion between the instances, we provide multiple views of the scene to cover its entire relevant extent unlike existing benchmarks with single-view inputs (see Fig. 1(a)).
@@
The proposed problem presents several challenges requiring the following key capabilities: i) achieving 3D understanding of the scene and view registration from multiple camera viewpoints without groundtruth 3D knowledge while handling occlusions, ii) aligning and comparing object instances without their pose information during training and testing, iii) learning representations that generalize to unseen object instances from seen or unseen categories at test time.
@@
To tackle these challenges, we introduce a novel method that takes as input multiple views of the same scene, projects them into a 3D voxel grid, generates a 3D object-centric representation for each instance, and predicts their labels by cross-correlating instances using an efficient attention mechanism. We utilize advances in differentiable rendering [26] and self-supervised learning [28] to supervise 3D representation learning. Specifically, we render the voxel representations for multiple viewpoints and match them with the given 2D views to produce geometrically consistent features. Additionally, we encourage our model to learn part-aware 3D representations by distilling features from 2D self-supervised model DINOv2 [28], enhancing the correspondence matching across instances and their alignment. Finally, since there is no prior benchmark for this task, we propose two new benchmarks – ToysAD-8K and PartsAD-15K, which include common objects categories as well as mechanical parts respectively to facilitate further research. Our method significantly outperforms baselines that do not compare instances with each other, and we rigorously analyze various aspects of our model and benchmarks.
@@ 2 Related Work
AD benchmarks. A key challenge in AD research is the scarcity of large datasets containing realistic anomalies. Earlier works [9, 31] focusing on high-level semantic anomalies often use existing classification datasets by treating a subset of classes as anomalies and the remainder as normal. There also exist several datasets containing real-world anomaly instances. For example, MVTec-AD [5] includes industrial objects with various defects like scratches, dents, and contaminations, Carrera et al. missing[8] presents various defects in nanofibrous material, and VisA [52] comprises complex industrial objects such as PCBs, as well as simpler objects like capsules and cashews, spanning a total of 12 categories. These datasets assume that objects are pose-aligned, and both normal images and their anomaly counterparts have the same pose. Zhou et al. missing[51] propose a pose-agnostic framework by introducing the PAD dataset, which comprises images of 20 LEGO bricks of animal toys from diverse viewpoints/poses. Unlike the works discussed above, we focus on AD in multi-object multi-view scene environments, where anomalies are predicted by assessing their mutual similarity with other objects in the scene. The proposed setting enables our framework to work seamlessly on novel object instances without requiring further training.
@@ 3 Method
Note that labels $y_{n}$ are defined relative to other objects in the scene. For example, consider a group of three coffee cups, of which two have yellow handles, and one has a blue handle. In this example, the latter cup is considered an anomaly.
@@ 4.1 Datasets
We have proposed two new benchmarks for multi-object AD: ToysAD-8K and PartsAD-15K. ToysAD-8K consists of real-world objects from multiple categories that enables to evaluate our model’s ability to generalize to unseen object categories. PartsAD-15K comprises a more diverse collection of mechanical object parts with arbitrary shapes, thus being free from any class-level inductive biases (e.g., a sheep with three legs are easy to be recognized as anomaly without comparison) of common objects. Both datasets include a wide range of fine-grained anomaly instances motivated by real-world applications in inspection and quality control. Scenes are generated with diverse backgrounds, illuminations, and camera viewpoints using photo-realistic ray tracing [38]. Next, we discuss the data generation for both datasets.
@@
To generate each scene, we randomly select a set of objects consisting of both the normal and anomalous variants of the same instance. Most scenes contain more normal objects than anomalies and few scenes have no anomalous instances. The resting poses for the objects are generated using rigid body simulation [4] by dropping the object to the floor. Objects are scaled and placed randomly in the scene while avoiding collisions We generated a total of $8K$ scenes. Each scene consists of $3$-$6$ objects rendered in $20$ views. For the training set, we randomly select $5K$ scenes from $39$ categories. We build two disjoint test sets. The first one (seen) contains $1K$ scenes from the seen categories but with unseen object instances. The second one (unseen) contains $2K$ scenes from the rest of the $12$ novel categories.
@@ 4.4 Ablations and Model Analysis
Limitations. Our benchmark and model have a few limitations. First, this work focuses on a specific set of anomalies common in manufacturing scenarios, potentially overlooking other types of anomalies found in different applications. Due to the high cost and difficulty of obtaining real damaged objects, our dataset primarily uses shapes of synthetic objects. Additionally, our model assumes that object instances are rigid and does not handle articulations or deformations. It also assumes objects are not neither touching nor fully occluded in the scene. Furthermore, when a scene contains mostly anomalies, the lack of ‘normal’ data points for comparison (except in cases like fractures or cracks) can make detection more challenging. Lastly, during real-world scene testing, factors such as noisy camera poses and the gap between synthetic and real environments may degrade the performance.  Fig. 9 shows some failure cases of our model.
@@ 5 Conclusion
In this paper, we have introduced a novel AD problem inspired by real-world applications along with two new benchmarks. The proposed task goes beyond the traditional AD setting and involves a cross study of objects in a scene from multiple camera viewpoints to identify the ‘odd-looking’ instances. We show that our model is robust to varying number of views and objects, and outperforms the baselines that do not consider cross-object correlations.

### v4

@@ Abstract
This paper introduces a novel anomaly detection (AD) problem aimed at identifying ‘odd-looking’ objects within a scene by comparing them to other objects present. Unlike traditional AD benchmarks with fixed anomaly criteria, our task detects anomalies specific to each scene by inferring a reference group of regular objects. To address occlusions, we use multiple views of each scene as input, construct 3D object-centric models for each instance from 2D views, enhancing these models with geometrically consistent part-aware representations. Anomalous objects are then detected through cross-instance comparison. We also introduce two new benchmarks, ToysAD-8K and PartsAD-15K as testbeds for future research in this task. We provide a comprehensive analysis of our method quantitatively and qualitatively on these benchmarks.
@@ 1 Introduction
Anomaly detection (AD) [11, 31] identifies patterns deviating from expected behavior. In standard vision AD benchmarks, anomalies arise from high-level factors such as presence of a previously unseen category [10, 33, 2] or low-level ones such as defects and variations in the normal object shape and texture [9, 6, 14]. These benchmarks typically assess objects in isolation, disregarding inter-object relationships and contextual dependencies. However, in many real-world scenarios, ‘normality’ is often context-dependent. For instance, in a production line for blue-handled coffee cups, a yellow handle is anomalous, and in a production line for yellow-handled coffee cups, a blue handle is anomalous (see Fig. 1(b)). Methods assuming a fixed standard of normality cannot accommodate such context-dependent anomalies.
Inspired by real-world visual inspection, where efficiency in automatically processing large volumes of the same product is crucial, we introduce a new AD problem with along two new benchmarks and a novel solution. Our setup (see Fig. 1(a)) consists of scenes containing multiple instances of the same object (e.g., coffee cup) including either all normal, or a mix of normal and anomalous instances. Each scene has at least two normal instances, which together serve as a scene-specific reference for ‘normality’. The objective is to detect anomalous instances while generalizing to previously unseen scenes including unseen objects and layouts. This approach facilitates cost-effective models that do not require retraining for new products. Our task diverges from traditional AD tasks in two key ways: (1) anomaly detection is scene-specific, requiring a comparative analysis of instances rather than solely relying on individual inspection – only certain anomalies (e.g., cracks and misaligned parts) are detectable at instance-level; (2) it involves multi-view scene analysis, capturing a comprehensive object model while addressing self-occlusion and inter-instance occlusion, unlike standard AD benchmarks that typically rely on single-view inputs (see Fig. 1(a)).
@@
The proposed problem presents several unique challenges requiring the following key capabilities: i) 3D-aware modeling: constructing accurate representations of object instances by capturing their geometry and appearance from multiple viewpoints while handling occlusions, ii) pose-agnostic comparison: aligning and comparing object instances effectively without their pose information during training and testing, iii) generalizable representations: learning representations that generalize to unseen object instances from both seen or unseen categories at test time.
@@
To tackle these challenges, we introduce a novel method that integrates multi-view scene analysis by projecting them into a 3D voxel grid and generating 3D object-centric representations. Our model classifies instances through cross-instance correlation using an efficient attention mechanism. Leveraging recent advances in differentiable rendering [28] and self-supervised learning [30], we ensure robust 3D representation learning by rendering voxel representations for multiple viewpoints and aligning them with corresponding 2D views for geometric consistency. Additionally, we enhance part-aware 3D representation learning by distilling features from 2D self-supervised model DINOv2 [30], improving correspondence matching and pose-invariant alignment. As no existing benchmarks address this task, we introduce two new benchmarks: ToysAD-8K including common objects categories, and PartsAD-15K focusing on mechanical parts, to foster further research. Our method significantly surpasses baselines that lack cross-instance comparison and part-awareness, and we provide an in-depth analysis of our model.
@@ 2 Related Work
AD benchmarks. A major challenge in AD research is the limited availability of large datasets with realistic anomalies. Earlier works [10, 33] focusing on high-level semantic anomalies often use existing classification datasets by treating a subset of classes as anomalies and the remainder as normal. There are also datasets containing real-world anomaly instances; for instance, MVTec-AD [5] includes industrial objects with various defects like scratches, dents, and contaminations, while Carrera et al. missing[9] presents various defects in nanofibrous material. The VisA [56] comprises complex industrial objects such as PCBs, as well as simpler objects like capsules and cashews, spanning a total of 12 categories. Zhou et al. missing[55] introduced a pose-agnostic framework by introducing the PAD dataset, which comprises images of 20 LEGO bricks of animal toys from diverse viewpoints/poses. Bhunia et al. missing[7] proposed a conditional AD detection technique by comparing a given 3D reference shape to a single object in a query image. In contrast, we target AD within multi-object, multi-view scene environments, where anomalies are defined relative to scenes and detected through cross-instance in the scene. Moreover, our method does not require 3D reference models which can be costly and tedious to obtain.
@@ 3 Method
Importantly $y_{n}$ is defined relative to other objects in the scene. For example, consider a group of three coffee cups, where two have yellow handles, and one has a blue handle. Here, the blue-handled cup is classified as an anomaly. Notably, we do not assume that normal instances outnumber anomalies; rather, we assume that anomalies exhibit instance-specific variations (e.g.,, fractures differing in location and lengths), while normal instances maintain high visual similarity.
@@ 4.1 Datasets
ToysAD-8K comprises real-world objects from multiple categories, enabling evaluation of our model’s generalization to unseen object categories. PartsAD-15K includes a diverse collection of mechanical parts with arbitrary shapes, eliminating class-level inductive biases present in common objects (e.g., a sheep with three legs are easy to be recognized as anomaly without comparison). Both datasets feature fine-grained anomaly instances inspired by real-world applications in inspection and quality control applications. Scenes are rendered with diverse backgrounds, illuminations, and camera viewpoints using photo-realistic ray tracing [40]. Next, we discuss the data generation for both datasets.
@@
Each scene is generated by randomly selecting a set of objects, including both the normal and anomalous variants of the same instance. While most scenes contain more normal objects than anomalies, some scenes includes only normal instances. Object resting poses are determined through rigid body simulation [4] where objects are dropped onto the floor. Objects are scaled and placed randomly in the scene while avoiding collisions We generated a total of $8K$ scenes. Each scene consists of $3$-$6$ objects rendered in $20$ views. For the training set, we randomly select $5K$ scenes from $39$ categories. We build two disjoint test sets. The first one (seen) contains $1K$ scenes from the seen categories but with unseen object instances. The second one (unseen) contains $2K$ scenes from the rest of the $12$ novel categories.
@@ 4.4 Ablations and Model Analysis
Limitations. Our benchmark and model come with a few limitations. First, this work targets specific anomalies common in manufacturing, which may limit its applicability to other types of anomalies found in different settings. Given the challenges and cost of obtaining real damaged objects, our dataset primarily uses shapes of synthetic objects. Additionally, our model assumes rigid objects and does not account for articulations or deformations. It also relies on objects being not touching nor not fully occluded in the scene. Furthermore, when a scene contains mostly anomalies, the lack of regular instances for comparison (except in cases like fractures or cracks) can further complicate the detection. Finally, real-world scene testing may introduce factors like noisy camera poses and the gap between synthetic and real environments, which could impact the performance. Fig. 9 illustrates some failure cases of our model. Most failure cases stem from obscure anomalies due to factors like low resolution, poor lighting, unresolved occlusions in cluttered scenes, limited multi-view images, or challenging objects (e.g., transparent textures).
@@ 5 Conclusion
We have introduced a novel AD problem inspired by real-world applications along with two new benchmarks. The proposed task goes beyond the traditional AD setting and involves a cross study of objects in a scene from multiple viewpoints to identify the ‘odd-looking’ instances. We show that our model is robust to varying number of views and objects, and outperforms the baselines that do not consider cross-object correlations and part-aware representations.
Acknowledgment. The authors acknowledge the support of Toyota Motor Europe.
Supplementary Material
This supplementary is structured as follows: we present additional details of the proposed datasets in  Appendix A, training details in  Appendix B, and additional results in  Appendix C.
Appendix A Data Generation Details
Our proposed scene AD datasets, ToysAD-8K and PartsAD-15K, are built upon two publicly available 3D shape datasets: Toys4K [38] (Creative Commons and royalty-free licenses) and ABC [24] (MIT license). For the ToysAD-8K, we selected 1,050 shapes from the Toys4K dataset, focusing on the most common real-world objects across 51 categories. A complete list of these categories is provided in Fig. A2, and the split of seen/unseen categories is shown in Tab. A1. On the other hand, PartsAD-15K is a non-categorical dataset. For this dataset, we randomly selected a subset of 4,200 shapes from the large-scale ABC. The test set of PartsAD-15K is comprised of unseen shapes. Given a 3D mesh model of an object, we automatically create anomalies by applying geometric deformations as described below.
Table A1: Dataset composition of ToysAD-8K
Categories
Seen dinosaur, fish, frog, monkey, light, lizard, orange, boat, dog, lion, pig, cookie, panda, chicken,
orange, ice, horse, car, airplane, cake, shark, donut, hat, cow, apple, bowl, hamburger, octopus,
giraffe, chess, bread, butterfly, cupcake, bunny, elephant, fox, deer, bus, bottle
Unseen mug, plate, robot, glass, sheep, shoe, train, banana, cup, key, penguin, hammer
Figure A1: Distribution of anomaly types in the proposed ToysAD-8K and PartsAD-15 datasets.
Figure A2: Distribution of object categories in ToysAD-8K dataset.
Figure A3: Anomaly detection performance (Accuracy) by anomaly type on the ToysAD-8K and PartsAD-15 datasets.
Different types of Anomalies. We consider the following anomaly types: cracks, fractures, geometric deformations (e.g., bumps, bends, and twists), misaligned parts (e.g., translation and rotation), material mismatch, and missing parts. Fig. A1 depicts the distribution of different anomaly types in our proposed dataset. Before applying deformations, we normalize the mesh vertices to $[-1,1]$. Then we synthetically generate realistic cracks and fractures following [36] to mimic the shape’s most geometrically natural breaking patterns. For geometric deformations, we use Blender’s [40] Simple Deform and Hook modifier to create global (e.g., bending and twisting) and local deformations (e.g., surface bumps), respectively. For translation, we randomly translate the vertices of a random part by an offset. The translation offset for each axis is randomly sampled from a uniform distribution of range $[0.04,0.08]$. For rotational anomalies, we apply a 3D rotational transformation to a random part. The rotation matrix is formed using a random rotation axis and a radian angle sampled from a uniform distribution of range $[0.2,0.4]$. The center of rotation is set to a fixed point at one of the connecting points between the anomalous part and the main body of the object. We also randomly change the texture of a part or region to create a material mismatch anomaly or completely remove a part to simulate a missing parts anomaly. Additionally, for PartsAD-15K, we use a different but geometrically similar instance in the same scene as an anomaly (referred to as ‘distinct’ in Fig. A1). To retrieve a geometrically similar shape, we use feature-based KNN clustering. Specifically, we extract DINOv2 features from multi-view images (rendered from multiple fixed viewpoints) of the 3D shapes, concatenate these features, and use the resulting concatenated features to build the KNN cluster. Then, we retrieve a similar 3D shape by querying the KNN cluster. To ensure the retrieved shape is geometrically similar, we calculate the Chamfer distance between the shapes and only accept a shape if the distance is below a certain threshold.
Scene construction. For generating scenes, we use resulting anomalies along with their normal instances. A scene can have more than one anomaly of the same instance, or it may not have any anomaly at all. For randomized object placement, we first determine each object’s resting pose by simulating a rigid body drop using Blender [40]. This simulation process is repeated for each object instance in the scene, including both normal and anomaly instances. Then, we place each object instance randomly using its obtained resting pose, while also ensuring that they do not collide with each other. We prevent collisions by maintaining a minimum distance between each pair of objects.
Quality checks. Our anomaly generation process is fully automated, with multiple quality checks to ensure reliable samples. For positional or rotational anomaly creation, we discard and regenerate any instance where a part detaches from the main body during deformation. Similarly, if removing a part results in an impractical shape, we try a different part instead. For fracture anomalies, we discard the sample where the fracture removes more than 90% or less than 10% of an object, regenerating a new fracture in these cases. Finally, we ensure the anomalous region of an object is visible from at least one viewpoint.
Assets. To generate a realistic scene environment, we use PBR materials [41] for floors and HDRI environment maps [53] for image-based lighting to illuminate scenes. We randomly select a pair of PBR material and HDRI environment maps from the assets to randomize the scene background. We employed Blender 2.93 [40] with Cycles ray-tracing renderer for photo-realistic rendering. Blender 2.93 is released under the GNU General Public License (GPL, or “free software”), and the PBR and HDRI maps are released under the CC0 license.
Rendering and view selection. We render each scene from various viewpoints sampled from a hemisphere around the scene origin. The viewpoint is parameterized in spherical coordinates where azimuth values are sampled uniformly over [0, $2\pi$) and elevation values are uniformly sampled in [$\pi/9$, $92\pi/$ 9]. We adjust the radius in the range [1.5, 2.5] to ensure that all the objects in the scene are visible (even if partially) in the rendered view.
Our framework can easily be trained with real-world manufacturing scene environments. Our model relies solely on 2D supervision, making the data collection process much easier. We also do not need precise annotation of 3D bounding boxes as they are not used during training. For annotating instance-wise anomaly labels, we can use 2D bounding boxes, which can be projected in 3D using Visual Hull [25], then used for coarse localization.
Appendix B Training Details
We train our model in two stages. In the first stage, we train it with just image and feature reconstruction losses. In the second stage, we train the model end-to-end with both reconstruction and binary classification losses. All the experiments are performed on a single NVIDIA A40 GPU with a batch size of 4, utilizing 28GB of GPU memory. The first stage takes 36 hours to complete, followed by an additional 24 hours for the second stage.
The ablation experiments are conducted on the same workstation with the same GPU by removing one or a few core components from the full method. Specifically, variant methods A and B take $36$ hours to train the first stage, while only taking $14$ hours to train the second stage. Regarding variant method C, it takes similar $36$ and $24$ hours for the two stages as in the full method.
We compare our method with Recons-Recog [35], ImVoxelNet [34], and DETR3D [46]. For the Recons-Recog approach, we use DGCNN [45] as a feature extractor. In terms of training time of these baseline methods, Recons-Recog takes $10$ hours to train, ImVoxelNet takes $24$ hours to train, and DETR3D takes $2$ days to converge.
Figure A4: We qualitatively compare our method with two baselines on the PartsAD-15K dataset. ImVoxelNet failed to detect the anomalous object in the first example. We use the red box for wrong predictions and the green box for correct predictions.
Appendix C Additional Results
Fig. A4 shows a qualitative comparison with two baseline methods: ImVoxelNet and DETR3D. Our method accurately predicts the anomalies, while the baselines perform poorly. In the first three examples, the baselines fail due to their inability to compare with other objects in the scene, which is necessary for correct predictions. In the last example, both DETR3D and our method predict correctly, while ImVoxelNet fails. In Fig. A5 and Fig. A6, we present additional qualitative results on ToysAD-8K and PartsAD-15K, respectively. We show a breakdown of accuracy across different anomaly types in Fig. A3 to interpret the model’s performance in various such scenarios.
Figure A5: Additional results on the unseen set of ToysAD-8K dataset. Each row shows multiple views of the input scene and the green box denotes our model’s prediction. For rows 1 to 3, the anomalies are easy to spot and self-explanatory. In row 4, there are two anomalies: one with broken outer parts at the back (see view 5), and another with a tilted roof (see views 1 and 2). For row 5, one leg is broken (see view 2). In row 6, the eye is missing (see view 5).
Figure A6: Additional results on the PartsAD-15K dataset. Each row shows multiple views of the input scene and the green box denotes our model’s prediction.